# A Randomized Exchange Algorithm for Optimal Design of Multi-Response Experiments

### v1

@@ 3.3.1 General Kiefer’s criteria
An important empirical observation is that during the execution of mREX a solution of (17) often corresponds to the full exchange of the weight of the design points with indices $i^{g}_{l}$ and $i^{s}_{k}$, i.e., the solutions are frequently $\alpha^{*}=\ell$ or $\alpha^{*}=u$. Therefore it is beneficial to rapidly determine whether one of these two cases holds. Concavity of $\phi$ yields that if $\phi^{\prime}(\ell\!\to\!u)\leq 0$ then $\alpha^{*}=\ell$ and if $\phi^{\prime}(u\!\to\!\ell)\leq 0$ then $\alpha^{*}=u$. If we are able to efficiently verify these conditions, many weight exchanges will be performed rapidly.
@@
    If neither of the above conditions is satisfied, we use a one-dimensional optimization procedure to find the maximum of $\phi$ on $[\ell+2\varepsilon,u-2\varepsilon]$. Similarly as above, for each $\alpha\in[\ell+2\varepsilon,u-2\varepsilon]$ the matrix $\mathbf{M}+\alpha\Delta$ is non-singular, therefore the optimization procedure does not encounter singularities. Moreover, $\phi$ is positive, concave and smooth at the optimization interval. Therefore, we can find the optimum $\alpha^{*}$ rapidly via a univariate optimization procedure (such as the base function optimize in R).
@@ 3.3.2 D-optimality
The complexity of the Faddeev-LeVerrier method applied directly to the $m\times m$ matrix $\mathbf{Q}$ is at least $O(m^{3})$ (e.g., [6]). Therefore, for larger $m$ the computational cost can be substantial. However, if $m$ is large, we typically have $2s<m$, and we can achieve significant numerical simplification by using the following approach. The generalized matrix determinant lemma (see Theorem 18.1.1 in [20]) implies for any $\alpha\in[\ell,u]$ and indices $i^{g}_{l}$, $i^{s}_{k}$
$\det\left[\mathbf{M}+\alpha\Delta\right]$ $=$ $\det\left[\mathbf{M}+\alpha(\mathbf{G}_{i^{g}_{l}}\mathbf{G}_{i^{s}_{k}}^{T}-\mathbf{G}_{\mathbf{k}_{k}}\mathbf{G}_{i^{s}_{k}}^{T})\right]$
$=$ $\det\left[\mathbf{M}+\alpha(\mathbf{G}_{i^{g}_{l}},\mathbf{G}_{i^{s}_{k}})(\mathbf{G}_{i^{g}_{l}},-\mathbf{G}_{i^{s}_{k}})^{T}\right]$
$=$ $\det\left[\mathbf{M}\right]\det\left[\mathbf{I}_{2s}+\alpha(\mathbf{G}_{i^{g}_{l}},-\mathbf{G}_{i^{s}_{k}})^{T}\mathbf{M}^{-1}(\mathbf{G}_{i^{g}_{l}},\mathbf{G}_{i^{s}_{k}})\right].$
@@
$\widetilde{\mathbf{\Delta}}=(\mathbf{G}_{i^{g}_{l}},-\mathbf{G}_{i^{s}_{k}})^{T}\mathbf{M}^{-1}(\mathbf{G}_{i^{g}_{l}},\mathbf{G}_{i^{s}_{k}})$
instead of $\mathbf{\Delta}$. This means that for the characteristic polynomial approach, we only need to apply the Faddeev-LeVerrier method to a $2s\times 2s$ matrix; in a typical application, e.g. with a binary response ($s=2$), the computation of characteristic polynomials for such small matrices is extremely rapid, independently of the number $m$ of parameters.
@@ 4.1 Variable response dimension and variable covariance matrices
where $\mathbf{y}_{\mathbf{x}}$ is the $k(\mathbf{x})$-dimensional response at the design point $\mathbf{x}\in\mathcal{X}$, $\boldsymbol{\beta}\in\mathbb{R}^{m}$ is the vector of parameters, and $\mathbf{F}(\mathbf{x})$ is a known matrix of size $m\times k(\mathbf{x})$. The vectors $\boldsymbol{\epsilon}_{\mathbf{x}}$ are assumed to be $\mathcal{N}_{k(\mathbf{x})}(\mathbf{0}_{k(\mathbf{x})},\sigma^{2}\boldsymbol{\Sigma}(\mathbf{x}))$ distributed and uncorrelated across different trials, with positive definite covariance matrices $\boldsymbol{\Sigma}(\mathbf{x})$ of dimension $k(\mathbf{x})\times k(\mathbf{x})$. In this model, the information matrix corresponding to a design $\mathbf{w}$ is
@@ 4.2 Nonlinear and generalized linear regression models
$\mathbf{y}_{\mathbf{x}}=\mathbf{F}^{T}(\mathbf{x},\boldsymbol{\beta})+\boldsymbol{\epsilon}_{\mathbf{x}},$ (21)
@@
$\textbf{G}(\mathbf{x}_{i},\boldsymbol{\beta}^{*})=\begin{bmatrix}\sqrt{v_{1}(\mathbf{x}_{i},\boldsymbol{\beta}^{*})}\mathbf{h}_{1}(\mathbf{x}_{i}),&\ldots&\sqrt{v_{s}(\mathbf{x}_{i},\boldsymbol{\beta}^{*})}\mathbf{h}_{s}(\mathbf{x}_{i})\end{bmatrix}\boldsymbol{\Sigma}^{-1/2},$
for known functions $v_{1},\ldots v_{s}$ that correspond to the distributions of the elements of $\mathbf{y}_{\mathbf{x}}$ and to the links $g_{1},\ldots,g_{s}$ (for some common functions $v$, see, e.g., [19]).
@@ 5 Application: dose-response studies
The primary objective of dose-response studies is to establish a balance between the therapeutic benefits of a drug and the potential risks associated with different dosage levels (see, e.g., [12], [26], [39]), that is, we measure a bivariate response corresponding to a level of efficacy and toxicity. However, the multi-response regression can be used to model any collection of possibly correlated measurements as a multivariate response to a given dose of a drug.
@@ 5.1 Bivariate Emax model
Local $D$-optimal minimally supported design for the Emax model (23) over the continuous interval $[a,b]$ is given by [36]. Here, “$D$-optimal minimally supported design” means that the design is $D$-optimal among all designs with the support size of $K=3$ points, which is the smallest support size of any non-singular design. Note that such designs need not be $D$-optimal among all designs: it may happen that the overall $D$-optimal design is supported on more than three points. The design given by [36] places weight $1/3$ on the following three points: $a$, $b$ and $x_{M}$, where
@@ Appendix B Proof of Theorem 2
Let $J\in\{1,\ldots,m\}$ be the number of passes through the repeat-until cycle of Algorithm 2. Let $\mathcal{S}_{0}:=\emptyset$ and $\mathbf{P}_{0}:=\mathbf{I}_{m}$. For $j=1,\ldots,J$, let $\mathbf{v}_{i}$ be the vector $\mathbf{v}$, $i^{*}_{j}$ be the index $i^{*}$, $\mathcal{S}_{j}$ be the set $\mathcal{S}$, $\widetilde{\mathbf{G}}_{j}$ be the matrix $\widetilde{\mathbf{G}}$, and $\mathbf{P}_{j}$ be the matrix $\mathbf{P}$, obtained in the $j$th pass of the repeat-until cycle (with $\mathbf{P}_{m}=\mathbf{0}_{m\times m}$ if $J=m$, i.e., if the break in Line 7 was realized).
@@
Because $\mathrm{tr}\left[\mathbf{P}_{j+1}\right]\leq\mathrm{tr}\left[\mathbf{P}_{j}\right]-1$ for each $j$, we have $\mathrm{tr}[\mathbf{P}_{j}]=0$ after at most $m$ steps (i.e., $\mathrm{tr}\left[\mathbf{P}_{J}\right]=0$ and $J\leq m$), which yields that $\mathbf{P}_{J}=\mathbf{0}_{m\times m}$. Consequently, $\mathcal{C}(\mathbf{M}(\mathcal{S}_{J}))=\mathcal{C}^{\perp}(\mathbf{P}_{J})=\mathbb{R}^{m}$, implying the non-singularity of the uniform design $\mathbf{w}$ on $\{\mathbf{x}_{i}:i\in\mathcal{S}_{J}\}=\{\mathbf{x}_{i}:i\in\mathcal{S}\}$. ∎

### v2

@@ 3.3.1 General Kiefer’s criteria
An important empirical observation is that during the execution of mREX a solution of (20) often corresponds to the full exchange of the weight of the design points with indices $i^{g}_{l}$ and $i^{s}_{k}$, i.e., the solutions are frequently $\alpha^{*}=\ell$ or $\alpha^{*}=u$. Therefore, it is beneficial to rapidly determine whether one of these two cases holds. Concavity of $\phi$ yields that if $\phi^{\prime}(\ell\!\to\!u)\leq 0$ then $\alpha^{*}=\ell$ and if $\phi^{\prime}(u\!\to\!\ell)\leq 0$ then $\alpha^{*}=u$. If we are able to efficiently verify these conditions, many weight exchanges will be performed rapidly.
@@
    If neither of the above conditions is satisfied, we use a one-dimensional optimization procedure to find the maximum of $\phi$ on $[\ell+2\varepsilon,u-2\varepsilon]$. Similarly as above, for each $\alpha\in[\ell+2\varepsilon,u-2\varepsilon]$ the matrix $\mathbf{M}+\alpha\Delta$ is nonsingular, therefore the optimization procedure does not encounter singularities. Moreover, $\phi$ is positive, concave and smooth at the optimization interval. Therefore, we can find the optimum $\alpha^{*}$ rapidly via a univariate optimization procedure (such as the base function optimize in R).
@@ 3.3.2 D-optimality
The complexity of the Faddeev-LeVerrier method applied directly to the $m\times m$ matrix $\mathbf{Q}$ is at least $O(m^{3})$ (e.g., [6]). Therefore, for larger $m$, the computational cost can be substantial. However, if $m$ is large, we typically have $2s<m$, and we can achieve significant numerical simplification by using the following approach. The generalized matrix determinant lemma (see Theorem 18.1.1 in [20]) implies for any $\alpha\in[\ell,u]$ and indices $i^{g}_{l}$, $i^{s}_{k}$
$\det\left[\mathbf{M}+\alpha\mathbf{\Delta}\right]$ $=$ $\det\left[\mathbf{M}+\alpha(\mathbf{G}(\mathbf{x}_{i^{g}_{l}})\mathbf{G}^{T}(\mathbf{x}_{i^{g}_{l}})-\mathbf{G}(\mathbf{x}_{i^{s}_{k}})\mathbf{G}^{T}(\mathbf{x}_{i^{s}_{k}})\right]$
$=$ $\det\left[\mathbf{M}+\alpha(\mathbf{G}(\mathbf{x}_{i^{g}_{l}}),\mathbf{G}(\mathbf{x}_{i^{s}_{k}}))(\mathbf{G}(\mathbf{x}_{i^{g}_{l}}),-\mathbf{G}(\mathbf{x}_{i^{s}_{k}}))^{T}\right]$
$=$ $\det\left[\mathbf{M}\right]\det\left[\mathbf{I}_{2s}+\alpha(\mathbf{G}(\mathbf{x}_{i^{g}_{l}}),-\mathbf{G}(\mathbf{x}_{i^{s}_{k}}))^{T}\mathbf{M}^{-1}(\mathbf{G}(\mathbf{x}_{i^{g}_{l}}),\mathbf{G}(\mathbf{x}_{i^{s}_{k}}))\right].$
@@
$\widetilde{\mathbf{\Delta}}=(\mathbf{G}(\mathbf{x}_{i^{g}_{l}}),-\mathbf{G}(\mathbf{x}_{i^{s}_{k}}))^{T}\mathbf{M}^{-1}(\mathbf{G}(\mathbf{x}_{i^{g}_{l}}),\mathbf{G}(\mathbf{x}_{i^{s}_{k}}))$
instead of $\mathbf{\Delta}$. This means that for the characteristic polynomial approach, we only need to apply the Faddeev-LeVerrier method to a $2s\times 2s$ matrix; in a typical application, e.g., with a binary response ($s=2$), the computation of characteristic polynomials for such small matrices is extremely rapid, independently of the number $m$ of parameters.
@@ 4.1 Variable response dimension and variable covariance matrices
where $\mathbf{y}_{\mathbf{x}}$ is the $s(\mathbf{x})$-dimensional response at the design point $\mathbf{x}\in\mathcal{X}$, $\boldsymbol{\beta}\in\mathbb{R}^{m}$ is the vector of parameters, and $\mathbf{F}(\mathbf{x})$ is a known matrix of size $m\times s(\mathbf{x})$. The vectors $\boldsymbol{\epsilon}_{\mathbf{x}}$ are assumed to be $\mathcal{N}_{s(\mathbf{x})}(\mathbf{0}_{s(\mathbf{x})},\sigma^{2}\boldsymbol{\Sigma}(\mathbf{x}))$ distributed and uncorrelated across different trials, with positive definite covariance matrices $\boldsymbol{\Sigma}(\mathbf{x})$ of dimension $s(\mathbf{x})\times s(\mathbf{x})$. In this model, the information matrix corresponding to a design $\mathbf{w}$ is
@@ 4.2 Nonlinear and generalized linear regression models
$\mathbf{y}_{\mathbf{x}}=\mathbf{F}(\mathbf{x},\boldsymbol{\beta})+\boldsymbol{\epsilon}_{\mathbf{x}},$ (24)
@@
$\textbf{G}(\mathbf{x}_{i},\boldsymbol{\beta}^{*})=\begin{bmatrix}\sqrt{v_{1}(\mathbf{x}_{i},\boldsymbol{\beta}^{*})}\mathbf{h}_{1}(\mathbf{x}_{i}),&\ldots,&\sqrt{v_{s}(\mathbf{x}_{i},\boldsymbol{\beta}^{*})}\mathbf{h}_{s}(\mathbf{x}_{i})\end{bmatrix}\boldsymbol{\Sigma}^{-1/2},$
for known functions $v_{1},\ldots,v_{s}$ that correspond to the distributions of the elements of $\mathbf{y}_{\mathbf{x}}$ and to the links $g_{1},\ldots,g_{s}$ (for some common functions $v$, see, e.g., [19]).
@@ 5 Application: dose-response studies
The primary objective of dose-response studies is to establish a balance between the therapeutic benefits of a drug and the potential risks associated with different dosage levels (see, e.g., [39], [12], [26]), that is, we often measure a bivariate response corresponding to a level of efficacy and toxicity. However, the multi-response regression can be used to model any collection of possibly correlated measurements as a multivariate response to a given dose of a drug, not only responses corresponding to efficacy and toxicity.
@@ 5.1 Bivariate Emax model
Local $D$-optimal minimally supported design for the Emax model (26) over the continuous interval $[a,b]$ is given by [36]. Here, “$D$-optimal minimally supported design” means that the design is $D$-optimal among all designs with the support size of $K=3$ points, which is the smallest support size of any nonsingular design. Note that such designs need not be $D$-optimal among all designs: it may happen that the overall $D$-optimal design is supported on more than three points. The design given by [36] places weight $1/3$ on the following three points: $a$, $b$ and $x_{M}$, where
@@ Appendix B Proof of Theorem 2
Let $J\in\{1,\ldots,m\}$ be the number of passes through the repeat-until cycle of Algorithm 3. Let $\mathcal{S}_{0}:=\emptyset$ and $\mathbf{P}_{0}:=\mathbf{I}_{m}$. For $j=1,\ldots,J$, let $\mathbf{v}_{j}$ be the vector $\mathbf{v}$, $i^{*}_{j}$ be the index $i^{*}$, $\mathcal{S}_{j}$ be the set $\mathcal{S}$, $\widetilde{\mathbf{G}}_{j}$ be the matrix $\widetilde{\mathbf{G}}$, and $\mathbf{P}_{j}$ be the matrix $\mathbf{P}$, obtained in the $j$th pass of the repeat-until cycle (with $\mathbf{P}_{m}=\mathbf{0}_{m\times m}$ if $J=m$, i.e., if the break in Line 7 was realized).
@@
Because $\mathrm{tr}\left[\mathbf{P}_{j+1}\right]\leq\mathrm{tr}\left[\mathbf{P}_{j}\right]-1$ for each $j$, we have $\mathrm{tr}[\mathbf{P}_{j}]=0$ after at most $m$ steps (i.e., $\mathrm{tr}\left[\mathbf{P}_{J}\right]=0$ and $J\leq m$), which yields that $\mathbf{P}_{J}=\mathbf{0}_{m\times m}$. Consequently, $\mathcal{C}(\mathbf{M}(\mathcal{S}_{J}))=\mathcal{C}^{\perp}(\mathbf{P}_{J})=\mathbb{R}^{m}$, implying the nonsingularity of the uniform design $\mathbf{w}$ on $\{\mathbf{x}_{i}:i\in\mathcal{S}_{J}\}=\{\mathbf{x}_{i}:i\in\mathcal{S}\}$. ∎